If a, b, c and d are the roots of the equation

2x⁴-4x³+6x²+10x+12=0.

Find the value of a²+b²+c²+d².


1
Expert's answer
2021-03-31T14:43:34-0400
"(a+b+c+d)^2"

"=(a+b)^2+2(a+b)(c+d)+(c+d)^2"

"=a^2+2ab+b^2+2ac+2ad+2bc+2bd"

"+c^2+2cd+d^2"

Then


"a^2+b^2+c^2+d^2"

"=(a+b+c+d)^2"

"-2(ab+ac+ad+bc+bd+cd)"

Use Vieta's formulas


"a+b+c+d=-\dfrac{-4}{2}=2"

"ab+ac+ad+bc+bd+cd=\dfrac{6}{2}=3"

Hence


"a^2+b^2+c^2+d^2=(2)^2-2(3)=-2"